# LiteGS: A High-performance Framework to Train 3DGS in Subminutes via System and Algorithm Codesign

### v2

@@ Introduction
Cluster-level Culling and Compaction
Based on the Morton-sorted array, we form clusters by simply grouping contiguous primitives (e.g., 128 primitives per cluster). We then compute an axis-aligned bounding box (AABB) for each cluster. Before rendering, we perform view frustum culling at the cluster level by testing the intersection of these AABBs with the camera’s view frustum, which allows us to cull large groups of invisible primitives at once. Subsequently, the data from all visible clusters are compacted into a contiguous memory buffer. This ensures that all subsequent computations benefit from coalesced memory access and reduces memory fragmentation.
Synergy with Sparse Gradient Updates
@@
SSIM↑/PSNR↑/LPIPS↓ Train↓ Num SSIM↑/PSNR↑/LPIPS↓ Train↓ Num SSIM↑/PSNR↑/LPIPS↓ Train↓ Num
@@
Table 1: Comparison of SSIM, PSNR, LPIPS, training time, and number of Gaussians across three datasets. * means the training iteration is 40K. For a fair comparison, LiteGS uses the same number of iterations as SSS. LiteGS-quality gets 0.780/24.72/0.133(better than others except SSS) in 30K iterations and takes 398s.
Opacity Gradient Variance Metric
The goal of densification is to adaptively add geometric detail in under-reconstructed regions. The original method primarily relies on the magnitude of the view-space position gradient to identify these regions, subsequently splitting or cloning the corresponding Gaussian primitives. However, this metric suffers from a fundamental ambiguity: a large gradient may not only signal a genuine under-reconstructed area that requires more primitives but could also represent a well-sampled but simply unconverged region.
This ambiguity is further exacerbated by the periodic opacity reset mechanism. Forcibly setting opacities to zero induces drastic, transient errors and gradients throughout the scene. Consequently, densification decisions made shortly after a reset are often unreliable, as it becomes difficult to distinguish whether a region has a true geometric defect or is merely in a transient state of optimization following the reset. To address this ambiguity, we introduce a new metric using the variance of the opacity gradient. Instead of merely measuring the magnitude of the gradient, we analyze its distribution across all pixels that observe the same Gaussian primitive. The intuition behind this is as follows:
This variance can be estimated efficiently. For each Gaussian primitive $i$, we compute the sum of squared gradients $S_{i}=\sum_{\pi\in\Pi}\sum_{(x,y)}(\frac{\partial\text{Loss}(I_{\pi}(x,y),\text{gt}_{\pi}(x,y))}{\partial o^{i}})^{2}$, the sum of gradients $M_{i}=\sum_{\pi\in\Pi}\sum_{(x,y)}\frac{\partial\text{Loss}(I_{\pi}(x,y),\text{gt}_{\pi}(x,y))}{\partial o^{i}}$, and the fragment count $C_{i}$. The densification metric is then calculated via $C_{i}\times\text{Var}=S_{i}-\frac{M_{i}^{2}}{C_{i}}$. Our proposed warp-based rasterization technique aggregates these pixel-level statistics with negligible overhead, enabling the efficient and low-cost computation of these moments.
Periodically resetting opacity can improve reconstruction quality by re-weighting all scene parameters, which are determined by the transmittance on each fragment. While performing more frequent resets is an intuitive idea, the disruptive impact of opacity resets adversely affects the densification process.
To mitigate these effects and allow for a higher reset frequency, we replace the hard reset with a gentler “opacity decay” mechanism. Instead of setting opacities directly to zero, we decay them by half. This approach ensures that the weights of all parameters in view are reset due to the increased transmittance, while also allowing the scene to recover to normal opacity levels much more quickly. Ultimately, with the opacity decay strategy, our densification hyperparameters are adjusted accordingly: the densification interval is extended to 5 epochs, which provides sufficient time for the scene to reach a “quasi-stable state” after an opacity adjustment for reliable variance metric calculation, while the opacity decay itself is performed every 10 epochs.
@@
SSIM↑/PSNR↑/LPIPS↓ Train↓ SSIM↑/PSNR↑/LPIPS↓ Train↓ SSIM↑/PSNR↑/LPIPS↓ Train↓
@@
Experiments
Datasets: We evaluated our method on three datasets: MipNeRF-360 (Barron et al. 2022), Tank & Temples (Knapitsch et al. 2017), and Deep Blending (Hedman et al. 2018). For MipNeRF-360 (Barron et al. 2022) dataset, we uniformly use the pre-downsampled images provided with the official dataset for all methods.
Baselines: We compare LiteGS against a comprehensive set of baselines, including foundational methods like the original 3DGS (Kerbl et al. 2023), high-quality reconstruction methods like 3DGS-MCMC (Kheradmand et al. 2024), SSS (Zhu et al. 2025), and recent acceleration-focused works such as TamingGS (Mallick et al. 2024), Mini-Splatting (Fang and Wang 2024a), and Mini-Splatting v2 (Fang and Wang 2024b).
Implementation Details: All experiments were performed on a single NVIDIA RTX 3090 GPU. We use PSNR, SSIM, and LPIPS as our evaluation metrics. For training, LiteGS adopts the same learning rates as TamingGS. Our densification starts after the third epoch, with splitting operations performed every five epochs. Opacity decay is applied every 10 epochs during the first 80% of training.
Parameter Scale: We carried out experiments on three distinct parameter scales. LiteGS-turbo runs at a small parameter scale and the number of training iterations (18K) is identical to Mini-splatting v2, aiming to showcase its high training performance. LiteGS-balance selects a moderate number of parameters to strike a balance between training accuracy and speed, using 30K training iterations. LiteGS-quality uses a parameter count comparable to the original 3DGS to achieve the highest possible reconstruction quality with also 30K training iterations.
Comparison with State-of-the-Art Methods
Table 1 presents a comprehensive comparison of LiteGS against other leading methods. Our framework demonstrates dominant and Pareto-optimal performance across all scales, achieving superior results in both rapid and high-quality reconstruction tasks.
Rapid Reconstruction. In the lightweight setting, LiteGS-turbo establishes a new SOTA for speed-quality trade-off. As shown in Table 1, compared to the original 3DGS, LiteGS-turbo reduces the number of primitives/parameters and the training time by up to 5.8x and 13.4x, respectively. It is also 1.4x faster than the currently fastest lightweight method, Mini-Splatting v2, with comparable accuracy (slightly better PSNR but worse LPIPS, as fewer points capture less high-frequency information). Another key advantage of LiteGS over Mini-splatting lies in the slow and predictable parameter expansion, but the latter uses an aggressive densify-and-prune strategy for lightweighting. Hence, LiteGS has a lower peak memory footprint, offering a significant advantage on devices with limited memory budgets.
High-Quality Reconstruction. In the million-scale parameter regime, LiteGS significantly outperforms its counterparts. LiteGS-balance, using only 1/3 to 1/2 the parameters of 3DGS-MCMC, is on par with its reconstruction quality but reducing the training time by an order of magnitude, i.e., from nearly an hour to 3-5 minutes. Even compared to the SOTA acceleration framework, TamingGS, LiteGS-balance cuts the training time by upto 3.7x (with an average of 2.97x), plus delivers significantly higher accuracy. Furthermore, LiteGS-quality sets a new SOTA for quality. Compared to 3DGS-MCMC, it not only speeds up training performance by 3.8x-7x (about 5-9 minutes), but also improves PSRN by 0.2-0.4 dB.
These results indicate that our high-performance framework, LiteGS, is able to deliver: 1) minute-level 3DGS training for the small parameter count scenarios (possibly second-level on new GPU like RTX 5090) without comprising reconstruction quality, 2) extremely high quality for the large parameter count tasks with remarkably less training overhead compared to virtually all existing solutions.
Metric Atomic Per-splat Tensor-core Warp-based
Scene: garden-5728k
Forward (ms) 17.82 12.58 6.81 3.01
Backward (ms) 51.84 27.31 9.28 6.93
Scene: garden-2000k
Forward (ms) 10.56 7.49 3.59 1.45
Backward (ms) 37.23 13.46 4.09 3.64
Scene: garden-500k
Forward (ms) 5.45 5.06 1.79 0.36
Backward (ms) 25.26 7.53 1.53 1.87
Table 3: Performance comparison of different rasterization methods on the ‘garden’ scene with varying numbers of primitives (5728k, 2000k, 500k). Our Warp-based Rasterizer shows a significant speedup in both forward and backward passes. Atomic Raster is the original implementaion in 3DGS. Per-splat Raster is the implementation from TamingGS which may be the most popular gaussian raster operator. Tensor-core is the recent SOTA works(Liao et al. 2025) and Warp-based Raster is our work.
Ablation Studies
This section validates the effectiveness of each key component of our framework through ablation studies.
Impact of Cluster-Cull-Compact. To isolate its effect, we disable our spatial clustering pipeline in the LiteGS-quality setting. The results in Table 2 show that it significantly increases training time (e.g., 36% slower, from 515s to 701s on Mip-NeRF 360) with only a minor impact on quality metrics. This reveals that our Cluster-Cull-Compact module effectively accelerates training.
@@
Impact of Warp-based Rasterizer. An end-to-end training time comparison is insufficient to fairly evaluate rasterizer performance, as our densification strategy may generate more visible primitives. Therefore, we list the cost breakdown of the operators in Table 3. Our warp-based rasterizer demonstrates a prominent speedup across all primitive scales, outperforming the original 3DGS atomic rasterizer by 7.0x-13.0x, popular per-splat rasterizers by  4.0x, and even recent SOTA methods (Liao et al. 2025) using Tensor Cores by  1.5x.
@@
Impact of Densification Strategy. Table 4 depicts the performance of each component in our densification strategy. In the setting without opacity decay, we use the hard opacity reset scheme and the same hyperparameters in the original 3DGS. In the configuration without the opacity gradient variance metric, per-pixel error (Bulò, Porzi, and Kontschieder 2024) is taken as the densification indicator. Using the original 3DGS densification hyperparameters and opacity reset policy could drop PSNR by 0.3-0.4 dB. This suggests that a more frequent reduction in overall scene opacity improves reconstruction quality, and our opacity decay mechanism avoids the instability of consecutive hard resets by allowing faster recovery. In the control experiment using per-pixel error, the PSNR drops by approx. 0.3-0.6 dB, indicating that the opacity gradient variance is a more robust metric than the error-based indicator.
Dataset LiteGS w/o Decay w/o Var w/o Both
Mip-NeRF 360
PSNR↑ 28.12 27.68 27.40 27.13
SSIM↑ 0.824 0.813 0.789 0.777
LPIPS↓ 0.209 0.222 0.235 0.259
Tanks & Temples
PSNR↑ 24.52 24.11 24.27 23.97
SSIM↑ 0.866 0.852 0.853 0.837
LPIPS↓ 0.145 0.166 0.163 0.190
Deep Blending
PSNR↑ 30.37 29.93 30.08 29.76
SSIM↑ 0.912 0.908 0.905 0.900
LPIPS↓ 0.233 0.239 0.245 0.259
@@
Table 4: Ablation study on our densification strategy. We compare our full model against variants without opacity decay, without the gradient variance metric (Var), or without both.
Conclusion

### v3

@@ 3.2.1 Morton Codes
3.2.2 Cluster-level Culling and Compaction
Based on the Morton-sorted array, we form clusters by simply grouping contiguous primitives (e.g., 128 primitives per cluster). We then compute an axis-aligned bounding box (AABB) for each cluster. Before rendering, we perform view frustum culling at the cluster level by testing the intersection of these AABBs with the camera’s view frustum, which allows us to cull large groups of invisible primitives at once. Subsequently, the data from all visible clusters are compacted into a contiguous memory buffer. This ensures that all subsequent computations benefit from coalesced memory access and reduces memory fragmentation. We provide a detailed explanation of the Morton code calculation and a visual illustration of the entire Cluster-Cull-Compact pipeline in Appendix A.3.
3.2.3 Synergy with Sparse Gradient Updates
@@
3.3 Opacity Gradient Variance Metric
The goal of densification is to adaptively add detail in under-reconstructed regions. The original method primarily relies on the magnitude of the view-space position gradient to identify these regions, subsequently splitting or cloning the corresponding Gaussian primitives. However, this metric suffers from a fundamental ambiguity: a large gradient may not only signal a genuine under-reconstructed area that requires more primitives but could also represent a well-sampled but simply unconverged region.
This ambiguity is exacerbated by the periodic opacity reset mechanism. Forcibly setting opacities to zero induces drastic, transient errors and gradients throughout the scene. Consequently, densification decisions made shortly after a reset are often unreliable, as it becomes difficult to distinguish whether a region has a true geometric defect or is merely in a transient state of optimization following the reset. To address this ambiguity, we introduce a new metric using the variance of the opacity gradient. Instead of merely measuring the magnitude of the gradient, we analyze its distribution across all pixels that observe the same Gaussian primitive. The intuition behind this is as follows:
This variance can be estimated efficiently. For each Gaussian primitive $i$, we compute the sum of squared gradients $S_{i}=\sum_{\pi\in\Pi}\sum_{(x,y)}(\frac{\partial\text{Loss}(I_{\pi}(x,y),\text{gt}_{\pi}(x,y))}{\partial o^{i}})^{2}$, the sum of gradients $M_{i}=\sum_{\pi\in\Pi}\sum_{(x,y)}\frac{\partial\text{Loss}(I_{\pi}(x,y),\text{gt}_{\pi}(x,y))}{\partial o^{i}}$, and the fragment count $C_{i}$. The densification metric is then calculated via $C_{i}\cdot\text{Var}=S_{i}-\frac{M_{i}^{2}}{C_{i}}$. Our proposed warp-based rasterization technique aggregates these pixel-level statistics with negligible overhead, enabling the efficient and low-cost computation of these moments.
Periodically resetting opacity can improve reconstruction quality by re-weighting all scene parameters, which are determined by the transmittance on each fragment. While performing more frequent resets is an intuitive idea, the disruptive impact of opacity resets adversely affects the densification process. To mitigate these effects and allow for a higher reset frequency, we replace the hard reset with a gentler “opacity decay” mechanism. Instead of setting opacities directly to zero, we decay them by half. This approach ensures that the weights of all parameters in view are reset due to the increased transmittance, while also allowing the scene to recover to normal opacity levels much more quickly. Ultimately, with the opacity decay strategy, our densification hyperparameters are adjusted accordingly: the densification interval is extended to 5 epochs, which provides sufficient time for the scene to reach a “quasi-stable state” after an opacity adjustment for reliable variance metric calculation, while the opacity decay itself is performed every 10 epochs.
4 Experiments
Datasets: We evaluated LiteGS on three datasets: Mip-NeRF 360 (Barron et al., 2022), Tank & Temples (Knapitsch et al., 2017), and Deep Blending (Hedman et al., 2018). For Mip-NeRF 360, we uniformly use the pre-downsampled images provided with the official dataset for all methods.
Baselines: We compare LiteGS against a comprehensive set of baselines, including foundational methods like the original 3DGS (Kerbl et al., 2023), high-quality reconstruction methods like 3DGS-MCMC (Kheradmand et al., 2024), SSS (Zhu et al., 2025), and recent acceleration-focused works such as TamingGS (Mallick et al., 2024), DashGaussian (Chen et al., 2025), Mini-Splatting (Fang & Wang, 2024a), and Mini-Splatting v2 (Fang & Wang, 2024b).
Implementation Details: All experiments are performed on a single NVIDIA RTX 3090 GPU. PSNR, SSIM, and LPIPS are used as evaluation metrics. LiteGS adopts the same learning rates as TamingGS. Our densification starts after the third epoch, with splitting operations performed every five epochs. Opacity decay is applied every $10$ epochs during the first 80% of training.
Parameter Scale: We carried out experiments on three distinct parameter scales. LiteGS-turbo runs at a small parameter scale and the number of training iterations (18K) is identical to Mini-splatting v2, aiming to showcase its high training performance. LiteGS-balance selects a moderate number of parameters to strike a balance between training accuracy and speed, using 30K training iterations. LiteGS-quality uses a parameter count comparable to the original 3DGS to achieve the highest possible reconstruction quality with also 30K training iterations. In addition to these three typical parameter scales, we provide more comparative results across a broader range of parameter scales in Section B.1.
4.1 Comparison with State-of-the-Art Methods
Table 1 presents the aggregate results from our comparative experiments with other leading methods, while the detailed per-scene metrics and comprehensive statistical analysis are provided in Section B.2. Qualitative visual comparisons demonstrating the reconstruction quality differences are available in Section B.3. Our framework demonstrates dominant and Pareto-optimal performance across all scales, achieving superior results in both rapid and high-quality reconstruction tasks. The metrics for 3DGS-MCMC and SSS in Table 1 differ from their published results due to their use of non-standard preprocessing on the Mip-NeRF 360 dataset, whereas Table 1 uses a unified, standard preprocessing for all methods. For the sake of a fair comparison, we also report results using their non-standard preprocessing pipeline in Section B.4.
Table 1: Comparison of SSIM, PSNR, LPIPS, training time, and number of Gaussians across three datasets. * means the training iteration is 40K. For a fair comparison, LiteGS uses the same number of iterations as SSS. LiteGS-quality gets 0.870/24.72/0.133 (better than others except SSS) in 30K iterations and takes 398s.
@@ 4.1 Comparison with State-of-the-Art Methods
SSIM/PSNR/LPIPS Train(s) Num SSIM/PSNR/LPIPS Train(s) Num SSIM/PSNR/LPIPS Train(s) Num
@@
Rapid Reconstruction. In the lightweight setting, LiteGS-turbo establishes a new SOTA for speed-quality trade-off. As shown in Table 1, compared to the original 3DGS, LiteGS-turbo reduces the number of primitives/parameters and the training time by up to 5.85.8x and 13.413.4x, respectively. It is also 1.51.5x faster than the currently fastest lightweight method, Mini-Splatting v2, with comparable accuracy (slightly better PSNR but worse LPIPS, as fewer points capture less high-frequency information). Another key advantage of LiteGS over Mini-splatting lies in the slow and predictable parameter expansion, but the latter uses an aggressive densify-and-prune strategy for lightweighting. Hence, LiteGS has a lower peak memory footprint, offering a significant advantage on devices with limited memory budgets.
High-Quality Reconstruction. In the million-scale parameter regime, LiteGS significantly outperforms its counterparts. LiteGS-balance, using only $1/3$ to $1/2$ the parameters of 3DGS-MCMC, is on par with its reconstruction quality but reducing the training time by an order of magnitude, i.e., from nearly an hour to 33-55 minutes. Even compared to the SOTA acceleration framework, TamingGS, LiteGS-balance cuts the training time by up to 3.73.7x (with an average of 33x), plus delivers significantly higher accuracy. DashGaussian [^2] Metrics are sourced from their GitHub, by reducing the number of primitives, shows a slight advantage in training speed over TamingGS, but suffers a significant drop in accuracy. Furthermore, LiteGS-quality sets a new SOTA for quality. Compared to 3DGS-MCMC, it not only speeds up training performance by 3.8x-7x (about 5-9 minutes), but also improves PSRN by 0.20.2-$0.4$ dB.
These results indicate that our high-performance framework, LiteGS, is able to deliver: 1) minute-level 3DGS training for the small parameter count scenarios without comprising reconstruction quality, 2) extremely high quality for the large parameter count tasks with remarkably less training overhead compared to virtually all existing solutions.
4.2 Ablation Studies
This section validates the effectiveness of each key component of LiteGS through ablation studies.
Impact of Cluster-Cull-Compact. To isolate its effect, we disable our spatial clustering pipeline in the LiteGS-quality setting. The results in Table 2 show that it significantly increases training time (e.g. 36% slower, from 515s to 701s on Mip-NeRF 360) with only a minor impact on quality metrics. This reveals that our Cluster-Cull-Compact module effectively accelerates training.
@@ 4.2 Ablation Studies
SSIM/PSNR/LPIPS Train(s) SSIM/PSNR/LPIPS Train(s) SSIM/PSNR/LPIPS Train(s)
@@
Impact of Warp-based Rasterizer. An end-to-end training time comparison is insufficient to fairly evaluate rasterizer performance, as our densification strategy may generate more visible primitives. Therefore, we list the cost breakdown of the operators in Table 3. Our warp-based rasterizer demonstrates a prominent speedup across all primitive scales, outperforming the original 3DGS atomic rasterizer by 7.0x-13.0x, popular per-splat rasterizers by  4.0x, and even recent SOTA methods (Liao et al., 2025) using Tensor Cores by  2x.
Table 3: Performance comparison of different rasterization methods on the ‘garden’ scene with varying numbers of primitives (500k, 2000k, 5728k). Our Warp-based Rasterizer shows a significant speedup in both forward and backward passes. Atomic Raster is the original implementaion in 3DGS. Per-splat Raster is the implementation from TamingGS which may be the most popular gaussian raster operator. Tensor-core GS (Liao et al., 2025) is the recent SOTA work for 3DGS inference and Warp-based Raster is our work.
Forward(ms) Backward(ms)
500k 2000k 5728k 500k 2000k 5728k
Atomic 5.45 10.56 17.82 25.26 37.23 51.84
Per-splat 5.06 7.49 12.58 7.53 13.46 27.31
Tensor-core 1.79 3.59 6.81 - - -
Warp-based 0.36 1.45 3.01 1.87 3.64 6.93
@@
Impact of Densification Strategy. Table 4 depicts the performance of each component in our densification strategy. In the setting without opacity decay, we use the hard opacity reset scheme and the same hyperparameters in the original 3DGS. In the setting without opacity gradient variance metric, per-pixel error (Bulò et al., 2024) is taken as the densification indicator. Using the original 3DGS densification hyperparameters and opacity reset policy could drop PSNR by 0.3-0.4 dB. This suggests that a more frequent reduction in overall scene opacity improves reconstruction quality, and our opacity decay mechanism avoids the instability of consecutive hard resets by allowing faster recovery. In the control experiment using per-pixel error, the PSNR drops by approx. 0.3-0.6 dB, indicating that the opacity gradient variance is a more robust metric than the error-based indicator.
@@
Table 4: Ablation study on our densification strategy. We compare our full model against variants without opacity decay (w/o Decay), without the gradient variance metric (w/o Var), and without both (w/o Both).
Mip-NeRF 360 Tanks & Temples Deep Blending
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
LiteGS 28.12 0.824 0.209 24.52 0.866 0.145 30.37 0.912 0.233
w/o Decay 27.68 0.813 0.222 24.11 0.852 0.166 29.93 0.908 0.239
w/o Var 27.40 0.813 0.235 24.27 0.853 0.163 30.08 0.905 0.245
w/o Both 27.13 0.777 0.259 23.97 0.837 0.190 29.76 0.900 0.259
5 Conclusion